OpenStudy (anonymous):

(7a)/(3a+3)-(5)/(4a-4)=(3a)/(2a+2)

OpenStudy (anonymous):

First need to get a common denominator.

OpenStudy (anonymous):

Just for reference, a=3 or -1/2

OpenStudy (anonymous):

What is the common denominator? And how do you solve it after you get the common denominator?

OpenStudy (anonymous):

The lcm should be (a+1)(a-1)

OpenStudy (anonymous):

Rewrite as \[\frac{ 7a }{ 3 }\left( \frac{ 1 }{a+1 } \right)-\frac{ 5 }{ 4 }\left( \frac{ 1 }{ a-1 } \right)=\frac{ 3 }{ 2 }\left( \frac{ 1 }{a+1 } \right)\]